40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Aug 4 23:03:56 2009, alberto, Update, Locking, IFO Alignment 
    Reply  Wed Aug 5 09:48:05 2009, alberto, Update, Locking, IFO Alignment 2009-08-09_FSStransPD.png2009-08-09_FSreflPD.png
       Reply  Thu Aug 6 09:33:08 2009, alberto, Update, Locking, FSS Transmitted and Reflected Power Trends 2009-08-06_PSL_trends200days.png2009-08-06_PSL_trends.png
          Reply  Thu Aug 6 10:32:45 2009, alberto, rob, Update, Locking, More PSL trends: NPRO, MOPA, FSS, PMC and MZ 2009-08-06_PSLtrends.png
Message ID: 1843     Entry time: Thu Aug 6 10:32:45 2009     In reply to: 1842
Author: alberto, rob 
Type: Update 
Category: Locking 
Subject: More PSL trends: NPRO, MOPA, FSS, PMC and MZ 

 Here we trended also the PMC and the MZ. The drop in the PMC happens at the same rate as the MOPA's.

That let us think that the FSS transmitteed power has gone down because of the reference cavity progressive misalignment to the laser beam.

We need to adjust that alignment sometime.

The drop in the NPRO output power (upper row, 3rd plot: Ch10 C1:PSL_126MOPA_126MON) accompained an increase of "fuzziness" in PMCTRANSPD and both coincided in time with the day we tempoarirly removed the flap from the laser chiller's chiller (July 14 2009).

Attachment 1: 2009-08-06_PSLtrends.png  21 kB  Uploaded Thu Aug 6 11:49:23 2009  | Hide | Hide all
2009-08-06_PSLtrends.png
ELOG V3.1.3-